# Organization Matters: A Qualitative Study of Organizational Dynamics in Red Teaming Practices for Generative AI

### v1

@@ Abstract.
The rapid integration of generative artificial intelligence (GenAI) across diverse fields underscores the critical need for red teaming efforts to proactively identify and mitigate associated risks. While previous research primarily addresses technical aspects, this paper highlights organizational factors that hinder the effectiveness of red teaming in real-world settings. Through qualitative analysis of 15 semi-structured interviews with red teamers from various organizations, we uncover challenges such as the marginalization of vulnerable red teamers, the invisibility of nuanced AI risks to vulnerable users until post-deployment, and a lack of user-centered red teaming approaches. These issues often arise from underlying organizational dynamics, including organizational resistance, organizational inertia, and organizational mediocracy. To mitigate these dynamics, we discuss the implications of user research for red teaming and the importance of embedding red teaming throughout the entire development cycle of GenAI systems.
@@
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: CSCW7article: 460publicationmonth: 11doi: 10.1145/3757641ccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
Red teaming for generative AI (GenAI) is a method for uncovering potential harms in AI products or systems by deliberately provoking failures through adversarial attacks or adversarial prompts (Inie et al., 2025). Framed as a preemptive measure to improve GenAI safety and protect end users, red teaming is widely supported among artificial general intelligence (AGI)  [^1] According to Schuett et al.  (Schuett et al., 2023), artificial general intelligence (AGI) refers to “AI systems that achieve or exceed human performance across a wide range of cognitive tasks”. AGI aims to develop general AI systems capable of generalizing beyond its original training objectives and executing tasks it was not explicitly designed for  (Raman et al., 2025). While AGI and generative AI (GenAI) are distinct—AGI referring to broad cognitive capabilities and GenAI emphasizes content generation—the terms can overlap in practice. researchers. In one survey, 98% of leading AGI experts somewhat or strongly agreed that AGI lab should conduct pre-deployment risk assessments, evaluate models for dangerous capabilities, commission third-party model audits, establish safety restrictions on model usage, and engage external red teams  (Schuett et al., 2023).
@@
As GenAI rapidly advances  (Bengio et al., 2024; AIS, 2024) and becomes integrated into various scenarios of daily life, red teaming efforts have lagged behind in addressing the growing risks posed to users and must now catch up with the expanding scope of generative AI applications  (Toner and Acharya, 2022; AIS, 2024). These applications now extend across a wide range of specialized contexts and user groups, including students learning human-computer interactions in higher education  (Kharrufa and Johnson, 2024; Bull and Kharrufa, 2024), K–12 students  (Bryant et al., 2020; Klopfer et al., 2024), medical professionals performing clinical tasks  (Melvin and Godwin, 2024; Manubolu and Peeriga, 2024; Umeton et al., 2023), and individuals undergoing psychological care  (Abrams, 2024), etc. These developments extend GenAI safety concerns beyond general-purpose chatbots (e.g., ChatGPT) and highlight the need to address risks emerging in situated, real-world contexts.
@@
In recent policy discussions and corporate initiatives, red teaming has emerged as a critical bulwark to safeguard users against the increasing risks associated with AI technologies  (Lucaj et al., 2023; Ganguli et al., 2022b, a). Streamlined red teaming have been adopted by organizations to ensure AI safety, including large technology companies (e.g. Microsoft  (Kumar, 2023), Google (Fabian, 2023), OpenAI (OpenAI et al., 2024)) and government agencies  (Lucaj et al., 2023; Ganguli et al., 2022b, a), often operating under the assumption that red teaming is one-size-fits-all solution to AI safety. However, limited attention has been paid to the harms that emerge from user behaviors in specific contexts (Dominique et al., 2024; Li et al., 2024; Ganguli et al., 2022b, a; Xu et al., 2021), the ways in which risks vary across user groups with heterogeneous values and ethical frameworks (Kirk et al., 2023; Feffer et al., 2024; Ganguli et al., 2022a), and the organizational practices of red teamers responsible for identifying and mitigating these risks.
@@
At the same time, organizational factors can impact, distort or reduce the effectiveness of red teaming (Ackerman and Clifford, 2021; Seid et al., 2025). Research on responsible AI has shown that such organizational dynamics shape how resources are allocated, how ethics are scoped and operationalized, and how the roles and identities of practitioners are constructed (Widder et al., 2024; Wang et al., 2023; Deng et al., 2023a). Yet, compared to technical aspects, the influence of organizational factors on red teaming for GenAI has received little attention in HCI and CSCW. In response, we adopt an organizational perspective to examine how red teaming is practiced across different organizations. We conducted interviews with 17 practitioners—including GenAI system red teamers, GenAI content moderators, GenAI product managers—across 17 organizations. Specifically, we probed from the following three research questions:
@@
This paper makes three key contributions. First, we provide an under-examined account of how red teaming workflows are enacted within real-world organizational settings, particularly those developing GenAI systems. Second, we offer an empirical analysis of the organizational dynamics that shape the effectiveness of red teaming implementation—specifically, organizational resistance, organizational inertia, and organizational mediocracy. Third, we discuss the potential of user research and the integration of red teaming throughout the GenAI development lifecycle as strategies for mitigating these organizaitonal challenges. Taken together, these contributions move beyond technical perspectives to offer a social, political, and organizational understanding of red teaming in the context of GenAI systems.
@@ 2. Related Work
As we study red teaming practices in organizations developing GenAI systems, we review prior research in two domains: red teaming practices, particularly in the context of GenAI and how Responsible AI (RAI) practitioners perceive and address AI-related harms in organizational settings. This review provides the foundation for our empirical analysis of how red teamers navigate these practices in real-world organizational contexts.
@@ 2.1. Red Teaming in the Context of Generative AI
The current body of research on red teaming methodologies has primarily focused on the generative model. Humphreys et al. identified four major vulnerabilities in GenAI: data poisoning, training data extraction, backdooring, and adversarial prompting  (Humphreys et al., 2024). Compared with red teaming for earlier AI systems, adversarial prompting is a distinct approach to GenAI and includes techniques such as prompt injection and adversarial examples  (Lar, 2024). To complement the limitations of manual red teaming, such as the inability to test every topic of interest  (Dominique et al., 2024), some companies have adopted automated tools, example databases, or prompt templates. Automated tools typically perform two key tasks: generating attack samples and conducting risk analysis  (Li et al., 2024). For example, Tencent’s Hunyuan model is a text-to-image diffusion transformer designed with fine-grained comprehension of both Chinese and English, designed to ensure compliance with various laws and regulatory attempts, such as the “Interim Measures for the Management of Generative AI Services”  (Li et al., 2024). Effective manual red teaming, meanwhile, requires experts to have both domain-specific knowledge and a deep understanding of how to probe AI models for unexpected behaviors  (Ganguli et al., 2022b). Prior research has also explored semi-automated techniques that incorporate human feedback as an intervention, aiming to combine the strengths and mitigate the limitations of both fully automated and fully manual red teaming approaches  (Xu et al., 2021).
Despite these methodological leaps, less attention has been given to defining user groups and assessing whether their values align with existing red-teaming methodologies. Additionally, issues such as ambiguous implementation processes and limited collaboration among domain experts, automated red-teaming tools, AI red teamers, and users remain unresolved. Our study explores these gaps by investigating how red teamers engage with these approaches in practice and how they contend with the related situational challenges that arise from their daily red teaming work.
@@ 2.2. The Heterogeneity of Perceived Harm Upon Users
When interactions from non-malicious users are taken into account, developing effective taxonomy and threat model to surface the risks before deployment becomes more complex due to the unpredictable and context-dependent nature of AI-generated outputs  (Mehrabi et al., 2023; Dorotic et al., 2024). Recognizing the limits of purely technical solutions, responsible AI (RAI) scholars increasingly emphasize the need to involve end users and domain stakeholders throughout the AI development lifecycle, particularly in defining fairness, safety, and harm. Smith et al.  (Smith et al., 2023) show that co-designing fairness metrics with stakeholders such as content providers on AI-driven platforms leads to metrics that better reflect lived experiences of unfairness and context-specific needs, bridging a gap often missed when developers act alone. Deng et al.  (Deng et al., 2022) found that industry practitioners struggled to apply one-size-fits-all “fairness toolkits” to their own AI projects. This suggests that these threat models need to be refined through direct participation of the practitioner to fit real-world use cases. Yildirim et al.  (Yildirim et al., 2023) further highlight how practitioners use human-AI guidelines such as the People + AI Guidebook not only to inform design decisions but also to educate colleagues and collaborate across teams, implying that practitioners’ feedback is essential in evolving these guidelines to address real-world challenges.
@@
Collectively, these findings suggest that proactive user and stakeholder engagement through co-design workshops and practitioner-centered tool development can help account for the heterogeneity in harmful perceptions and better align GenAI systems with the values and expectations of diverse users. At the same time, fully realizing participatory RAI also requires supporting the marginalized workers who carry out much of this work. Hence, embracing these participatory approaches in GenAI development and providing institutional support for the people doing this invisible labor is a promising strategy to anticipate and mitigate a broader range of harms. This paper examines how these heterogeneous perceptions of harm manifest in actual red teaming practices within organizations. While previous work has focused on developing theoretical frameworks for user involvement, we empirically investigate how red teamers navigate these varying definitions of harm in practical settings with insufficient guidance or user research.
@@ 2.3. Red Teaming as Organizational Practices
Organizational factors critically shape the effectiveness of red teaming and broader responsible AI (RAI) efforts. As researchers benchmark GenAI systems against human-value criteria through risk audits, safety scorecards, and ethics analyses  (O’Neil et al., 2024; Jabbour and Janapa Reddi, 2024; Iyer, 2024), evaluation results and metrics may still fail to capture qualitative error patterns or address critical conceptual flaws. These bechmark performances can also vary significantly depending on evaluation context and prompting strategies  (Roy et al., 2024; Jin et al., 2024; Tao et al., 2024; Madaio et al., 2022). These limitations reflect a pattern of means-ends decoupling, in which well-intentioned practices fall short of achieving their intended outcomes  (Bromley and and, 2012). As organizational theory suggests, when RAI practices are difficult to evaluate and the relationship between means and ends is opaque, organizations tend to develop increasingly complex internal structures, persist in perpetual reform cycles, and divert resources away from core goals  (Bromley and and, 2012; Ali et al., 2023). For red teaming, this means efforts may become procedural formalities rather than effective safeguards, with resources spent on process compliance instead of uncovering and addressing real risks. Organizations also face strong counter-incentives to disclose red teaming results due to potential reputational risks  (Ganguli et al., 2022a), creating policy-practice decoupling where organizations may adopt AI safety policies without implementing corresponding practices  (Bromley and and, 2012). RAI researchers have observed similar patterns of decoupling in organizational ethics efforts, where policies, practices, and outcomes diverge, and collective action within organizations is often absent  (Ali et al., 2023). Ali et al.  (Ali et al., 2023) identify three core challenges to re-couple policies, practices, and outcomes in organizational RAI practices: (1) prioritizing ethics under product launch pressures; (2) quantifying ethics in metrics-focused environments; and (3) team reorganizations that disrupt institutional memory and key relationships. These dynamics often individualize risk, making the act of raising ethical concerns a personal liability—particularly for workers from marginalized backgrounds.
@@ 3.2. Interview Protocol
We developed our semi-structured interview protocol iteratively through the first two interviews (P1, P2). Following these interviews, we reviewed and refined some questions for better clarity, and their sequence was optimized to enhance conversational flow. The interviews for 15 participants who took part in the formal study lasted an average of 46 minutes, with a range of 26 to 102 minutes. The same protocol was used for both industry and academic red teamers to ensure comparability across participant responses. For example, we first asked each participant to introduce themselves and describe their experience with generative AI (e.g., “How long have you been using generative AI, and what are your overall impressions of it?”). Participants then described their specific roles, responsibilities, and typical workflows (e.g., “Could you briefly describe your position and your typical day at work?”). Next, we focused on their understanding and practical approaches to red teaming (e.g., “How would you describe red teaming work? Can you explain your strategies for testing and mitigating harmful AI-generated content?”). We explored their experiences and the challenges encountered while doing the red teaming related works (e.g., “What types of harmful content do you typically test for, and what challenges have you faced?”). Participants also discussed their emotional reactions to encountering harmful content (e.g., “How do you typically respond to harmful AI-generated content? Could you provide examples?”). We further explored the use and perceptions of automated tools and databases (e.g., “Do you use any automated tools or databases to assist with your red teaming tasks? How helpful are they compared to human team members?”) and the workplace dynamics (Cheon et al., 2022) within their organizations (e.g., “How do you work with colleagues or departments inside and outside your organization before and after identifying risks?”). Finally, we addressed stress management and workplace support mechanisms (e.g., “How does your organization protect your mental health when exposed to harmful content? What practices are employed by leadership to support your well-being?”). Participants were also encouraged to envision and describe their ideal red teaming work environment and workflow.

### v2

@@ Abstract.
The rapid integration of generative artificial intelligence (GenAI) across diverse fields underscores the critical need for red teaming efforts to proactively identify and mitigate associated risks. While previous research primarily addresses technical aspects, this paper highlights organizational factors that hinder the effectiveness of red teaming in real-world settings. Through qualitative analysis of 17 semi-structured interviews with red teamers from various organizations, we uncover challenges such as the marginalization of vulnerable red teamers, the invisibility of nuanced AI risks to vulnerable users until post-deployment, and the lack of user-centered red teaming approaches. These issues often arise from underlying organizational dynamics, including organizational resistance, organizational inertia, and organizational mediocracy. To mitigate these dynamics, we discuss the implications of user research for red teaming and the importance of embedding red teaming throughout the entire development cycle of GenAI systems.
@@
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW460publicationmonth: 11doi: 10.1145/3757641ccs: Human-centered computing Collaborative and social computing
@@ 1. Introduction
Red teaming for generative AI (GenAI) is a method for uncovering potential harms in AI products or systems by deliberately provoking failures through adversarial attacks or adversarial prompts  (Inie et al., 2025). Framed as a preemptive measure to improve GenAI safety and protect end users, red teaming is widely supported among artificial general intelligence (AGI)  [^1] According to Schuett et al.  (Schuett et al., 2023), artificial general intelligence (AGI) refers to “AI systems that achieve or exceed human performance across a wide range of cognitive tasks.” AGI aims to develop general AI systems capable of generalizing beyond its original training objectives and executing tasks it was not explicitly designed for  (Raman et al., 2025). While AGI and generative AI (GenAI) are distinct—AGI referring to broad cognitive capabilities and GenAI emphasizes content generation—the terms can overlap in practice.researchers. In one survey, 98% of leading AGI experts somewhat or strongly agreed that AGI labs should conduct pre-deployment risk assessments, evaluate models for dangerous capabilities, commission third-party model audits, establish safety restrictions on model usage, and engage external red teams  (Schuett et al., 2023).
@@
As GenAI rapidly advances  (Bengio et al., 2024; AIS, 2024) and becomes integrated into various scenarios of daily life, red teaming efforts have lagged behind in addressing the growing risks posed to users and must now catch up with the expanding scope of generative AI applications  (Toner and Acharya, 2022; AIS, 2024). These applications now extend across a wide range of specialized contexts and user groups, including students learning human-computer interactions in higher education  (Kharrufa and Johnson, 2024; Bull and Kharrufa, 2024), K–12 students  (Bryant et al., 2020; Klopfer et al., 2024), medical professionals performing clinical tasks  (Melvin and Godwin, 2024; Manubolu and Peeriga, 2024; Umeton et al., 2023), and individuals undergoing psychological care  (Abrams, 2024), etc. These developments extend GenAI safety concerns beyond general-purpose chatbots (e.g., ChatGPT) and highlight the need to address risks emerging in situated and user-centered contexts.
@@
In recent policy discussions and corporate initiatives, red teaming has emerged as a critical bulwark to safeguard users against the increasing risks associated with AI technologies  (Lucaj et al., 2023; Ganguli et al., 2022b, a). Streamlined red teaming has been adopted by organizations to ensure AI safety, including large technology companies (e.g., Microsoft  (Kumar, 2023), Google (Fabian, 2023), OpenAI (OpenAI et al., 2024)) and government agencies  (Lucaj et al., 2023; Ganguli et al., 2022b, a), often operating under the assumption that red teaming is a one-size-fits-all solution to AI safety. However, limited attention has been paid to the harms that emerge from user behaviors in specific contexts (Dominique et al., 2024; Li et al., 2024; Ganguli et al., 2022b, a; Xu et al., 2021), the ways in which risks vary across user groups with heterogeneous values and ethical frameworks (Kirk et al., 2023; Feffer et al., 2024; Ganguli et al., 2022a), and the organizational practices of red teamers responsible for identifying and mitigating these risks.
@@
At the same time, organizational dynamics can impact, distort or reduce the effectiveness of red teaming (Ackerman and Clifford, 2021; Seid et al., 2025). Research on responsible AI has shown that such organizational dynamics shape how resources are allocated, how ethics are scoped and operationalized, and how the roles and identities of practitioners are constructed (Widder et al., 2024; Wang et al., 2023; Deng et al., 2023a). Yet, compared to technical aspects, the influence of organizational dynamics on red teaming for GenAI has received little attention in HCI and CSCW. In response, we adopt an organizational perspective to examine how red teaming is practiced across different organizations. We conducted interviews with 17 practitioners—including GenAI system red teamers, GenAI content moderators, and GenAI product managers—across various organizations. Specifically, we probed the following three research questions:
@@
This paper makes three key contributions. First, we provide an underexamined account of how red teaming workflows are enacted within real-world organizational settings, particularly those developing GenAI systems. Second, we offer an empirical analysis of the organizational dynamics that shape the effectiveness of red teaming implementation—specifically, organizational resistance, organizational inertia, and organizational mediocracy. Third, we discuss how user research and the integration of red teaming throughout the GenAI development lifecycle can help mitigate these organizational challenges. Taken together, these contributions move beyond technical perspectives to offer a social, political, and organizational understanding of red teaming in the context of GenAI systems.
@@ 2. Related Work
As we study red teaming practices in organizations developing GenAI systems, we review prior research in two domains: red teaming practices, particularly in the context of GenAI and how Responsible AI (RAI) practitioners perceive and address AI-related harms within organizational settings. This review provides the foundation for our empirical analysis of how red teamers navigate these practices in real-world organizational contexts.
@@ 2.1. Red Teaming in the Context of Generative AI
The current body of research on red teaming methodologies has primarily focused on the generative models. Humphreys et al. identified four major vulnerabilities in GenAI: data poisoning, training data extraction, backdooring, and adversarial prompting  (Humphreys et al., 2024). Compared with red teaming for earlier AI systems, adversarial prompting is a distinct approach to GenAI and includes techniques such as prompt injection and adversarial examples  (Lar, 2024). To complement the limitations of manual red teaming, such as the inability to test every topic of interest  (Dominique et al., 2024), some companies have adopted automated tools, example databases, or prompt templates. Automated tools typically perform two key tasks: generating attack samples and conducting risk analysis  (Li et al., 2024). For example, Tencent’s Hunyuan model is a text-to-image diffusion transformer designed with fine-grained comprehension of both Chinese and English, designed to ensure compliance with various laws and regulatory attempts, such as the “Interim Measures for the Management of Generative AI Services”  (Li et al., 2024). Effective manual red teaming, meanwhile, requires experts to have both domain-specific knowledge and a deep understanding of how to probe AI models for unexpected behaviors  (Ganguli et al., 2022b). Prior research has also explored semi-automated techniques that incorporate human feedback as an intervention, aiming to combine the strengths and mitigate the limitations of both fully automated and fully manual red teaming approaches  (Xu et al., 2021).
Despite these methodological leaps, less attention has been given to defining user groups and assessing whether their values align with existing red teaming methodologies. Additionally, issues such as ambiguous implementation processes and limited collaboration among domain experts, automated red teaming tools, AI red teamers, and users remain unresolved. Our study explores these gaps by investigating how red teamers engage with these approaches in practice and how they contend with the related situational challenges that arise from their daily red teaming work.
@@ 2.2. The Heterogeneity of Perceived Harm Upon Users
When interactions from non-malicious users are taken into account, developing effective taxonomies and threat models to surface the risks before deployment becomes more complex due to the unpredictable and context-dependent nature of AI-generated outputs  (Mehrabi et al., 2023; Dorotic et al., 2024). Recognizing the limits of purely technical solutions, responsible AI (RAI) scholars increasingly emphasize the need to involve end users and domain stakeholders throughout the AI development lifecycle, particularly in defining fairness, safety, and harm. Smith et al.  (Smith et al., 2023) show that co-designing fairness metrics with stakeholders such as content providers on AI-driven platforms leads to metrics that better reflect lived experiences of unfairness and context-specific needs, bridging a gap often missed when developers act alone. Deng et al.  (Deng et al., 2022) found that industry practitioners struggled to apply one-size-fits-all “fairness toolkits” to their own AI projects. This suggests that these threat models need to be refined through direct participation of practitioners to fit real-world use cases. Yildirim et al.  (Yildirim et al., 2023) further highlight how practitioners use human-AI guidelines such as the People + AI Guidebook not only to inform design decisions but also to educate colleagues and collaborate across teams, implying that practitioners’ feedback is essential in evolving these guidelines to address real-world challenges.
@@
Collectively, these findings suggest that proactive user and stakeholder engagement through co-design workshops and practitioner-centered tool development can help account for the heterogeneity in perceptions of harm and better align GenAI systems with the values and expectations of diverse users. At the same time, fully realizing participatory RAI also requires supporting the marginalized workers who carry out much of this work. Hence, embracing these participatory approaches in GenAI development and providing institutional support for the people doing this invisible labor is a promising strategy to anticipate and mitigate a broader range of harms. This paper examines how these heterogeneous perceptions of harm manifest in actual red teaming practices within organizations. While previous work has focused on developing theoretical frameworks for user involvement, we empirically investigate how red teamers navigate these varying definitions of harm in practical settings with insufficient guidance or user research.
@@ 2.3. Red Teaming as Organizational Practices
Organizational factors critically shape the effectiveness of red teaming and broader responsible AI (RAI) efforts. As researchers benchmark GenAI systems against human-value criteria through risk audits, safety scorecards, and ethics analyses  (O’Neil et al., 2024; Jabbour and Janapa Reddi, 2024; Iyer, 2024), evaluation results and metrics may still fail to capture qualitative error patterns or address critical conceptual flaws. These benchmark performances can also vary significantly depending on evaluation context and prompting strategies  (Roy et al., 2024; Jin et al., 2024; Tao et al., 2024; Madaio et al., 2022). These limitations reflect a pattern of means-ends decoupling, in which well-intentioned practices fall short of achieving their intended outcomes  (Bromley and and, 2012). As organizational theory suggests, when RAI practices are difficult to evaluate and the relationship between means and ends is opaque, organizations tend to develop increasingly complex internal structures, persist in perpetual reform cycles, and divert resources away from core goals  (Bromley and and, 2012; Ali et al., 2023). For red teaming, this means efforts may become procedural formalities rather than effective safeguards, with resources spent on process compliance instead of uncovering and addressing real risks. Organizations also face strong counter-incentives to disclose red teaming results due to potential reputational risks  (Ganguli et al., 2022a), creating policy-practice decoupling where organizations may adopt AI safety policies without implementing corresponding practices  (Bromley and and, 2012). RAI researchers have observed similar patterns of decoupling in organizational ethics efforts, where policies, practices, and outcomes diverge, and collective action within organizations is often absent  (Ali et al., 2023). Ali et al.  (Ali et al., 2023) identify three core challenges to re-couple policies, practices, and outcomes in organizational RAI practices: (1) prioritizing ethics under product launch pressures; (2) quantifying ethics in metrics-focused environments; and (3) team reorganizations that disrupt institutional memory and key relationships. These dynamics often individualize risk, making the act of raising ethical concerns a personal liability—particularly for workers from marginalized backgrounds.
@@ 3.2. Interview Protocol
We developed our semi-structured interview protocol iteratively through the first two interviews (P1, P2). Following these interviews, we reviewed and refined some questions for better clarity, and their sequence was optimized to enhance conversational flow. The interviews for the remaining 15 participants who took part in the formal study lasted an average of 46 minutes, with a range of 26 to 102 minutes. The same protocol was used for both industry and academic red teamers to ensure comparability across participant responses. For example, we first asked each participant to introduce themselves and describe their experience with generative AI (e.g., “How long have you been using generative AI, and what are your overall impressions of it?”). Participants then described their specific roles, responsibilities, and typical workflows (e.g., “Could you briefly describe your position and your typical day at work?”). Next, we focused on their understanding and practical approaches to red teaming (e.g., “How would you describe red teaming work? Can you explain your strategies for testing and mitigating harmful AI-generated content?”). We explored their experiences and the challenges encountered while doing the red teaming related works (e.g., “What types of harmful content do you typically test for, and what challenges have you faced?”). Participants also discussed their emotional reactions to encountering harmful content (e.g., “How do you typically respond to harmful AI-generated content? Could you provide examples?”). We further explored the use and perceptions of automated tools and databases (e.g., “Do you use any automated tools or databases to assist with your red teaming tasks? How helpful are they compared to human team members?”) and the workplace dynamics (Cheon et al., 2022; Cheon and Erickson, 2025) within their organizations (e.g., “How do you work with colleagues or departments inside and outside your organization before and after identifying risks?”). Finally, we addressed stress management and workplace support mechanisms (e.g., “How does your organization protect your mental health when exposed to harmful content? What practices are employed by leadership to support your well-being?”). Participants were also encouraged to envision and describe their ideal red teaming work environment and workflow.